# ‘Do as I say not as I do’: A Semi-Automated Approach for Jailbreak Prompt Attack against Multimodal LLMs

### v2

@@
Abstract.
Large Language Models (LLMs) have seen widespread applications across various domains due to their growing ability to process diverse types of input data, including text, audio, image and video. While LLMs have demonstrated outstanding performance in understanding and generating contexts for different scenarios, they are vulnerable to prompt-based attacks, which are mostly via text input. In this paper, we introduce the first voice-based jailbreak attack against multimodal LLMs, termed as Flanking Attack, which can process different types of input simultaneously towards the multimodal LLMs. Our work is motivated by recent advancements in monolingual voice-driven large language models, which have introduced new attack surfaces beyond traditional text-based vulnerabilities for LLMs. To investigate these risks, we examine the state-of-the-art multimodal LLMs, which can be accessed via different types of inputs such as audio input, focusing on how adversarial prompts can bypass its defense mechanisms. We propose a novel strategy, in which the disallowed prompt is flanked by benign, narrative-driven prompts. It is integrated in the Flanking Attack which attempts to humanizes the interaction context and execute the attack through a fictional setting. Further, to better evaluate the attack performance, we present a semi-automated self-assessment framework for policy violation detection. We demonstrate that Flank Attack is capable of manipulating state-of-the-art LLMs into generating misaligned and forbidden outputs, which achieves an average attack success rate ranging from 0.67 to 0.93 across seven forbidden scenarios. These findings highlight both the potency of prompt-based obfuscation in voice-enabled contexts and the limitations of current LLMs’ moderation safeguards. With a comprehensive evaluation of Flank Attack, this work establishes a replicable testing framework for adversarial robustness evaluation in multimodal LLMs. It highlights the urgent need for advanced defense strategies to address the challenges posed by evolving, context-rich attacks.
@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
1. Introduction
@@
The rapid advancement of Large Language Models (LLMs) has fostered significant progress across various domains, from natural language processing to multimodal interactions involving audio, text, and images. However, as these models become increasingly integral to critical applications, concerns regarding their robustness and susceptibility to adversarial attacks have garnered particular attention. One type of adversarial attacks, known as the jailbreak attack, aims to circumvent the internal constraints and safeguards, thereby obtaining the prohibited contents and response from LLMs (Greshake et al., 2023; Wei et al., 2024). It has since become a significant threat to LLMs, mostly focusing on the craft of deliberate text prompt to mislead the LLMs (Xu et al., 2024).
@@
In recent years, there has been a growing body of research works investigating the limitations of LLMs in handling adversarial inputs across multiple languages and modalities (Shayegani et al., 2023b). While numerous constraints have been implemented to actively defense such threats, the jailbreak techniques to bypass such defense continue to evolve (Shen et al., 2024a; Yu et al., 2024; Chang et al., 2024). Most of these studies emphasize text-based or multilingual environments for LLMs, resulting in the curation of jailbreak prompts, as illustrated in Figure. 1.
Figure 1. Example prompt and completions for refusals on disallowed categories.
[…]
[¡short description¿]¡long description¿
@@
As such, we have seen collective efforts for systematic and comprehensive benchmark of different jailbreak attacks (Chu et al., 2024; Yu et al., 2024). On the one hand, many works aim at identifying the underlying strategies of existing jailbreak prompts and their effectiveness for a more systematic understanding. In Shen et al. (Shen et al., 2024a) and Yu et al. (Yu et al., 2024), the jailbreak prompts from various online sources are analyzed to deliberately violate OpenAI’s policies. On the other hand, new techniques for creating novel and effective jailbreak prompts, which target different aspects of LLMs continue to emerge. These include methods such as multi-language mixture attacks leveraging low-resource languges by Upadhayay et al. (Upadhayay and Behzadan, 2024) and system prompt leaking attack by Hui et al. (Hui et al., 2024).
Figure 2. Overview of Voice Jailbreak(Chu et al., 2024)
[…]
[¡short description¿]¡long description¿
@@
In latest developments, multimodal LLMs have demonstrated exceptional capabilities in processing diverse input types, such as audio, vision, and text. For example, GPT-4o is designed to handle direct audio input with a user-friendly interface (Hurst et al., 2024), while Gemini features strong generalist capabilities across modalities, along with advanced understanding and reasoning performance in various domains (Team et al., 2023). This has inevitably introduces novel attack surfaces for security concerns (Liu et al., 2024), which have been highlighted in most recent works for voice mode of GPT-4o (Shen et al., 2024b) and video-based multimodal LLMs (Huang et al., 2025). While these works have made valuable contributions, the extent and potential harm caused by jailbreak prompts to multimodal LLMs remain unclear. Among these, audio-based jailbreak attacks pose a unique challenge, as they exploit the auditory input capabilities of LLMs, potentially circumventing established defense mechanisms. As shown in Figure. 2, with dedicated design of jailbreak prompt, the attacker can lead the multimodal LLMs for a misaligned and potentially forbidden response to the query, even such forbidden question is explicitly listed as part of the user policy for LLMs usage.
@@
Furthermore, in this paper, we present a novel, simple and universal black-box jailbreak attack method, named Flanking Attack, to generate highly effective prompts for state-of-the-art multimodal LLMs. We have developed a systematic semi-automated framework for audio-based jailbreak prompt attacks. With the framework, we are able to provide a more comprehensive and insightful evaluation for the robustness of mulitimodal LLMs’ defenses strategies. Although some studies present different security threats for a range of modalities (Liu et al., 2024), different languages (Upadhayay and Behzadan, 2024) even system information (Hui et al., 2024), our work demonstrates the specific challenges faced by audio-enabled LLMs. We observe that, for these LLMs, they may not benefit much from language diversity in training process while exhibiting distinct vulnerabilities. Appendix D delves deeper into the limitations of the multimodal LLMs, providing a nuanced analysis of their performance and highlighting potential areas for future improvements. To this end, our work contributes to a better understanding of LLMs’ defensive capabilities, offering a foundation for developing more resilient LLMs that can withstand sophisticated audio-based attacks in the future.
@@
Contributions. Our contributions are listed as follows.
We perform a systematic benchmarking of audio-based jailbreak attacks against the state-of-the-art multimodal LLMs. This work presents a thorough approach specifically designed to exploit potential vulnerabilities within monolingual, audio-enabled LLMs. Upon investigation, we observe that current approaches for defending multimodal LLMs from jailbreak prompt attacks are generally effective when addressing queries consisting solely of pure forbidden questions.
We propose a novel, simple and universal audio-based jailbreak attack framework for LLMs. By leveraging prompt-based guidance and franking the jailbreak query with benign ones, we propose Franking Attack, which can successfully bypass the defense mechanism in the multimodal LLMs. In total, we have evaluated 2,100 prompts across seven forbidden scenarios, yielding an impressive ASR result of 0.81. This finding highlights the importance of prompt formulation in jailbreak attack strategy, as it can drastically enhance the likelihood of bypassing defensive mechanisms. The framework’s adaptability and scalability make it a valuable tool for future studies on audio-based jailbreak attacks against multimodal LLMs, providing a structured approach to assess and replicate adversarial scenarios across similar models.
We develop a semi-automated approach to evaluate jailbreak prompt attacks against multimodal LLMs. To further enhance the generation and assessment of jailbreak prompt attacks for multimodal LLMs, a new semi-automated approach is introduced. Unlike traditional method, we feature an aligned multimodal LLM to facilitate the evaluation of responses for policy compliance. We anticipate this approach can be adapted for future research of multimodal LLMs so that they will progressively improve response quality and better align with safety guidelines.
2. Background
2.1. Large Language Models
@@
LLMs, such as OpenAi’s GPT family, and LLaMa, are a type of generative artificial intelligence designed to comprehend and generate human-like outputs. These models utilize deep learning architecture, particularly Transformers, to interpret and generate coherent and contextual text. As noted by (Pandurengan, [n. d.]), ‘Generative AI, often referred to as GenAI, involves systems that have the capacity to generate content autonomously, and large language models (LLMs) are a prominent embodiment of this concept’.
@@
Recent works have demonstrated LLMs’ tremendous potential for achiveing human-like intelligence (Achiam et al., 2023; Touvron et al., 2023; Chen et al., 2019), benefiting from large-scale training datasets along with a significant number of model parameters. However, concerns have been raised regarding the safety and security of LLMs, particularly their misuse by malicious actors. These risks involve a wide range of issues, including social engineering and data breaches (Glukhov et al., 2023). To mitigate such risks, various methods have been presented to regulate LLMs usage, including fine-tuning LLMs to enhance ethical standards compliance (Achiam et al., 2023; Touvron et al., 2023) and using third-party censorship systems to identify and filter inappropriate inputs or outputs (Glukhov et al., 2023). Despite these efforts, existing defenses have been empirically circumvented (Chang et al., 2024; Chu et al., 2024; Shayegani et al., 2023b; Huang et al., 2023; Russinovich et al., 2024). It highlights the critical need for further research to ensure the safe and ethical udage of LLMs.
2.2. LLMs safety and its limitation
@@
LLMs may fabricate facts (referred to as ‘hallucinations’), create divisive content, or reproduce prejudice, hate speech or stereotypes (Glukhov et al., 2023). These problems largely arise from the large and diverse data sets use for pre-training. Reinforcement Learning with Human Feedback (RLHF) is thus considered as mitigation to align LLMs with human values to address these issues (Glukhov et al., 2023; Bai et al., 2022).
@@
According to GenAi, Gemini has a decreased propensity to produce damaging information or have hallucinations(Team et al., 2023). Undesirable behaviors may arise when instructions to labelers are inadequately specified during the reward model data collection phase of the Reinforcement Learning with Human Feedback (RLHF) pipeline(Achiam et al., 2023). The process involves injecting policy-driven guidelines, sometimes referred to as ’constitutions,’ which guide the model in handling sensitive topics such as political neutrality. For example, in topics like elections, the model is trained to avoid taking sides, ensuring that responses maintain a neutral point of view. This method is inspired by Constitutional AI, where human feedback is used to revise responses and select safer outputs.(Team et al., 2023)
@@
Taking Gemini as an example, the zero-sample reasoning ability (Kojima et al., 2022) to modify answers and choose between multiple candidate answers. For example, in topics like elections, the model is trained to avoid taking sides, ensuring that responses maintain a neutral point of view. This method is inspired by Constitutional AI, where human feedback is used to revise responses and select safer outputs.
@@
According to recent research on supervised fine-tuning (SFT), balancing the harmlessness and helpfulness of language model responses remains a significant challenge. While a response such as, ‘I cannot help with that because it violates X policy’, refer to Appendix D, ensures safety by avoiding harmful content, it often fails to meet user expectations for helpfulness. The difficulty lies in striking an appropriate balance between these two objectives within the model’s fine-tuning framework (Team et al., 2023). Furthermore, the landscape of harmful query patterns is highly dynamic, which complicates fast mitigation. Models must quickly adapt to newly discovered harmful query patterns while also generalizing their defenses to address a wide array of harmful inputs. To address this, advanced chain-of-thought reasoning based on safety policy concepts has been introduced, allowing models to operate at a higher level of abstraction rather than responding solely at the fine-grained example level (Zhang et al., 2023a).
@@
There are limitations to LLM safety and alignment, particularly regarding the ability of LLMs to ‘self-censor’ and consistently ensure their outputs are always permissible. Theoretically, it has been established that the expected results of LLMs can be arbitrarily misaligned as long as the model can produce misaligned outputs with a non-zero probability, regardless of how the outputs are aligned (as measured by a specific metric) (Xu et al., 2023a). Furthermore, because LLMs must store information to function effectively, it was acknowledged that they could not maintain privacy or avoid data poisoning without making significant performance trade-offs (El-Mhamdi et al., 2022).
@@
Figure 3. Responses to a malicious instruction by the LLAMA2-7B-CHAT model under different generation configurations.(Huang et al., 2023)
[…]
[¡short description¿]¡long description¿
3. Related Works
3.1. Adversarial Attack
@@
Machine learning algorithms are known to be vulnerable to adversarial attacks, in which the carefully crafted inputs can result in producing consistently erroneous outputs (Goodfellow et al., 2014). Understanding adversarial attacks in the context of an LLM presents several challenges. LLMs are complicated since they are large-scale, imaginative, situational, multi-modal, and increasingly intertwined into intricate ecosystems (e.g. large language model based autonomous agents (Shayegani et al., 2023b)). As such, the threat presented by adversarial attackers exhibits distinct behaviors that require meticulous examination. Thus, accurate threat models is important to guide the creation of principled defenses. Here we list some motivational examples for adversarial attack on an LLM:
  1. (1)
@@
Figure 4. A taxonomy of concepts covered in the survey. (Shayegani et al., 2023b)
[…]
[¡short description¿]¡long description¿
3.2. Multi-modal Attack
The target model of a multi-modal attack accepts input from multiple modalities (such as text, images, audio, etc.)(Girdhar et al., 2023). These attacks exploit the increased complexity and introduction of new vulnerabilities by combining different types of input data. Additional modalities open up fresh attack vectors. Traditional textual alignment methods frequently fail to protect the joint embedding space that these multi-modal inputs generate.
Figure 5. Adversarial Embedding Space Attack(Russinovich et al., 2024)
[…]
[¡short description¿]¡long description¿
3.3. Jailbreak Prompt Attacks

### v3

@@
Abstract
Large Language Models (LLMs) have seen widespread applications across various domains due to their growing ability to process diverse types of input data, including text, audio, image and video. While LLMs have demonstrated outstanding performance in understanding and generating contexts for different scenarios, they are vulnerable to prompt-based attacks, which are mostly via text input. In this paper, we introduce the first voice-based jailbreak attack against multimodal LLMs, termed as Flanking Attack, which can process different types of input simultaneously towards the multimodal LLMs. Our work is motivated by recent advancements in monolingual voice-driven large language models, which have introduced new attack surfaces beyond traditional text-based vulnerabilities for LLMs. To investigate these risks, we examine the state-of-the-art multimodal LLMs, which can be accessed via different types of inputs such as audio input, focusing on how adversarial prompts can bypass its defense mechanisms. We propose a novel strategy, in which the disallowed prompt is flanked by benign, narrative-driven prompts. It is integrated in the Flanking Attack which attempts to humanizes the interaction context and execute the attack through a fictional setting. Further, to better evaluate the attack performance, we present a semi-automated self-assessment framework for policy violation detection. We demonstrate that Flanking Attack is capable of manipulating state-of-the-art LLMs into generating misaligned and forbidden outputs, which achieves an average attack success rate ranging from 0.67 to 0.93 across seven forbidden scenarios. These findings highlight both the potency of prompt-based obfuscation in voice-enabled contexts and the limitations of current LLMs’ moderation safeguards. With a comprehensive evaluation of Flank Attack, this work establishes a replicable testing framework for adversarial robustness evaluation in multimodal LLMs. It highlights the urgent need for advanced defense strategies to address the challenges posed by evolving, context-rich attacks.
@@ Abstract
I Introduction
@@
The rapid advancement of Large Language Models (LLMs) has fostered significant progress across various domains, from natural language processing to multimodal interactions involving audio, text, and images. However, as these models become increasingly integral to critical applications, concerns regarding their robustness and susceptibility to adversarial attacks have garnered particular attention. One type of adversarial attacks, known as the jailbreak attack, aims to circumvent the internal constraints and safeguards, thereby obtaining the prohibited contents and response from LLMs [1, 2]. It has since become a significant threat to LLMs, mostly focusing on the craft of deliberate text prompt to mislead the LLMs [3].
@@
In recent years, there has been a growing body of research works investigating the limitations of LLMs in handling adversarial inputs across multiple languages and modalities [4]. While numerous constraints have been implemented to actively defense such threats, the jailbreak techniques to bypass such defense continue to evolve [5, 6, 7]. Most of these studies emphasize text-based or multilingual environments for LLMs, resulting in the curation of jailbreak prompts, as illustrated in Figure. 1.
Figure 1: Example prompt and completions for refusals on disallowed categories.
@@
As such, we have seen collective efforts for systematic and comprehensive benchmark of different jailbreak attacks [8, 6]. On the one hand, many works aim at identifying the underlying strategies of existing jailbreak prompts and their effectiveness for a more systematic understanding. In Shen et al. [5] and Yu et al. [6], the jailbreak prompts from various online sources are analyzed to deliberately violate OpenAI’s policies. On the other hand, new techniques for creating novel and effective jailbreak prompts, which target different aspects of LLMs continue to emerge. These include methods such as multi-language mixture attacks leveraging low-resource languges by Upadhayay et al. [9] and system prompt leaking attack by Hui et al. [10].
Figure 2: Overview of Voice Jailbreak[8]
@@
In latest developments, multimodal LLMs have demonstrated exceptional capabilities in processing diverse input types, such as audio, vision, and text. For example, GPT-4o is designed to handle direct audio input with a user-friendly interface [11], while Gemini features strong generalist capabilities across modalities, along with advanced understanding and reasoning performance in various domains [12]. This has inevitably introduces novel attack surfaces for security concerns [13], which have been highlighted in most recent works for voice mode of GPT-4o [14] and video-based multimodal LLMs [15]. While these works have made valuable contributions, the extent and potential harm caused by jailbreak prompts to multimodal LLMs remain unclear. Among these, audio-based jailbreak attacks pose a unique challenge, as they exploit the auditory input capabilities of LLMs, potentially circumventing established defense mechanisms. As shown in Figure. 2, with dedicated design of jailbreak prompt, the attacker can lead the multimodal LLMs for a misaligned and potentially forbidden response to the query, even such forbidden question is explicitly listed as part of the user policy for LLMs usage.
@@
Furthermore, in this paper, we present a novel, simple and universal black-box jailbreak attack method, named Flanking Attack, to generate highly effective prompts for state-of-the-art multimodal LLMs. We have developed a systematic semi-automated framework for audio-based jailbreak prompt attacks. With the framework, we are able to provide a more comprehensive and insightful evaluation for the robustness of mulitimodal LLMs’ defenses strategies. Although some studies present different security threats for a range of modalities [13], different languages [9] even system information [10], our work demonstrates the specific challenges faced by audio-enabled LLMs. We observe that, for these LLMs, they may not benefit much from language diversity in training process while exhibiting distinct vulnerabilities. Appendix D delves deeper into the limitations of the multimodal LLMs, providing a nuanced analysis of their performance and highlighting potential areas for future improvements. To this end, our work contributes to a better understanding of LLMs’ defensive capabilities, offering a foundation for developing more resilient LLMs that can withstand sophisticated audio-based attacks in the future.
@@
Contributions. Our contributions are as follows.
A systematic benchmarking of audio-based jailbreak attacks on state-of-the-art multimodal LLMs. This work presents the first comprehensive evaluation of audio-based jailbreak attacks targeting monolingual, audio-enabled multimodal LLMs. Our study reveals that existing defense approaches are generally effective when addressing queries containing isolated, explicitly forbidden questions.
A novel and universal audio-based jailbreak attack framework. We propose the Franking Attack, a simple yet powerful jailbreak framework that strategically franks the jailbreak query between benign ones to evade detection and leverages the prompt-based guidance. We significantly increase ASR, yielding an impressive 0.81 ASR across 2,100 prompts ranging across seven forbidden scenarios. This finding highlights the importance of prompt formulation in jailbreak attack strategy, as it can drastically enhance the likelihood of bypassing defensive mechanisms. The framework’s adaptability and scalability make it a valuable tool for future studies on audio-based jailbreak attacks against multimodal LLMs, providing a structured approach to assess and replicate adversarial scenarios across similar models.
A semi-automated approach to evaluate jailbreak attacks against multimodal LLMs. To further enhance the generation and evaluation of jailbreak attacks for multimodal LLMs, we introduce a new semi-automated approach that incorporate an aligned multimodal LLM to assess policy compliance of model responses. Unlike traditional method, our approach facilitates scalable and consistent evaluations. We anticipate this approach will support future research on aligning multimodal LLMs with safety protocols by enabling more efficient and reliable benchmarking.
II Background
II-A Large Language Models
@@
LLMs, such as OpenAi’s GPT family, and LLaMa, are a type of generative artificial intelligence designed to comprehend and generate human-like outputs. These models utilize deep learning architecture, particularly Transformers, to interpret and generate coherent and contextual text. As noted by [16], ‘Generative AI, often referred to as GenAI, involves systems that have the capacity to generate content autonomously, and large language models (LLMs) are a prominent embodiment of this concept’.
@@
Recent works have demonstrated LLMs’ tremendous potential for achiveing human-like intelligence [17, 18, 19], benefiting from large-scale training datasets along with a significant number of model parameters. However, concerns have been raised regarding the safety and security of LLMs, particularly their misuse by malicious actors. These risks involve a wide range of issues, including social engineering and data breaches [20]. To mitigate such risks, various methods have been presented to regulate LLMs usage, including fine-tuning LLMs to enhance ethical standards compliance [17, 18] and using third-party censorship systems to identify and filter inappropriate inputs or outputs [20]. Despite these efforts, existing defenses have been empirically circumvented [7, 8, 4, 21, 22]. It highlights the critical need for further research to ensure the safe and ethical udage of LLMs.
II-B LLMs safety and its limitation
@@
LLMs may fabricate facts (referred to as ‘hallucinations’), create divisive content, or reproduce prejudice, hate speech or stereotypes [20]. These problems largely arise from the large and diverse data sets use for pre-training. Reinforcement Learning with Human Feedback (RLHF) is thus considered as mitigation to align LLMs with human values to address these issues [20, 23].
@@
According to GenAi, Gemini has a decreased propensity to produce damaging information or have hallucinations[12]. Undesirable behaviors may arise when instructions to labelers are inadequately specified during the reward model data collection phase of the Reinforcement Learning with Human Feedback (RLHF) pipeline[17]. The process involves injecting policy-driven guidelines, sometimes referred to as ’constitutions,’ which guide the model in handling sensitive topics such as political neutrality. For example, in topics like elections, the model is trained to avoid taking sides, ensuring that responses maintain a neutral point of view. This method is inspired by Constitutional AI, where human feedback is used to revise responses and select safer outputs.[12]
@@
Taking Gemini as an example, the zero-sample reasoning ability [24] to modify answers and choose between multiple candidate answers. For example, in topics like elections, the model is trained to avoid taking sides, ensuring that responses maintain a neutral point of view. This method is inspired by Constitutional AI, where human feedback is used to revise responses and select safer outputs.
@@
According to recent research on supervised fine-tuning (SFT), balancing the harmlessness and helpfulness of language model responses remains a significant challenge. While a response such as, ‘I cannot help with that because it violates X policy’, refer to Appendix D, ensures safety by avoiding harmful content, it often fails to meet user expectations for helpfulness. The difficulty lies in striking an appropriate balance between these two objectives within the model’s fine-tuning framework [12]. Furthermore, the landscape of harmful query patterns is highly dynamic, which complicates fast mitigation. Models must quickly adapt to newly discovered harmful query patterns while also generalizing their defenses to address a wide array of harmful inputs. To address this, advanced chain-of-thought reasoning based on safety policy concepts has been introduced, allowing models to operate at a higher level of abstraction rather than responding solely at the fine-grained example level [25].
@@
There are limitations to LLM safety and alignment, particularly regarding the ability of LLMs to ‘self-censor’ and consistently ensure their outputs are always permissible. Theoretically, it has been established that the expected results of LLMs can be arbitrarily misaligned as long as the model can produce misaligned outputs with a non-zero probability, regardless of how the outputs are aligned (as measured by a specific metric) [26]. Furthermore, because LLMs must store information to function effectively, it was acknowledged that they could not maintain privacy or avoid data poisoning without making significant performance trade-offs [27].
@@
Figure 3: Responses to a malicious instruction by the LLAMA2-7B-CHAT model under different generation configurations.[21]
III Related Works
III-A Adversarial Attack
@@
Machine learning algorithms are known to be vulnerable to adversarial attacks, in which the carefully crafted inputs can result in producing consistently erroneous outputs [28]. Understanding adversarial attacks in the context of an LLM presents several challenges. LLMs are complicated since they are large-scale, imaginative, situational, multi-modal, and increasingly intertwined into intricate ecosystems (e.g. large language model based autonomous agents [4]). As such, the threat presented by adversarial attackers exhibits distinct behaviors that require meticulous examination. Thus, accurate threat models is important to guide the creation of principled defenses. Here we list some motivational examples for adversarial attack on an LLM:
  1. 1.
@@
Figure 4: A taxonomy of concepts covered in the survey. [4]
III-B Multi-modal Attack
The target model of a multi-modal attack accepts input from multiple modalities (such as text, images, audio, etc.)[29]. These attacks exploit the increased complexity and introduction of new vulnerabilities by combining different types of input data. Additional modalities open up fresh attack vectors. Traditional textual alignment methods frequently fail to protect the joint embedding space that these multi-modal inputs generate.
Figure 5: Adversarial Embedding Space Attack[22]
III-C Jailbreak Prompt Attacks